Given the initial-value problem ?′′ + 3?′ + 2? = 4?, ?(0) = 3, ?′(0) =...

  1. Given the initial-value problem ?′′ + 3?′ + 2? = 4?, ?(0) = 3, ?′(0) = 1,

  2. Find its homogeneous solution using the Constant Coefficient approach (10pts)

  3. Find is particular solution using the Annihilator method. (10pts)

  4. Find the general solution that satisfies the initial conditions. (5pts)
Answer #1


963, 400021 y + 3y +2y = un herrn D.E, ylla phy W. ki, 4-04, egy 2 m bry +3 by + 2 y = un (2 ² + 30+ 2) y a un comparing wi mtizo mitt = 0 Im m = -2, ma-1 Complementary ffon utinis Yeni = C, e24 + Gen To find particular solution: p (n) = Axts y 024yni= -2C, 82 _ CE +2 py no use frol=1 9102- - 2 C, 60 e + 2 1 = -20 -12 +1 - 1 = -29-l2 - 2ata solve or 6 9 = 2,+ch 1-24